You decide to travel to a star 62 light-years from Earth at a speed that tells you the distance is only 25 lightyears. How many years would it take you to make the trip?



1
Expert's answer
2021-05-13T09:20:24-0400

The equation for the contracted distance:


"L=L_0\sqrt{1-\frac{v^2}{c^2}}."

Express the velocity "v":


"v=c\sqrt{1-\frac{L^2}{L_0^2}}."

The time required to pass distance "L" at speed "v":


"t=\frac Lv=\frac{L}{c\sqrt{1-(L/L_0)^2}}=861.545·10^6\text{ s},\\\space\\
t=27^\text y 3^\text m26^\text d."